OpenStudy (anonymous):

An isosceles triangle has a base of m and sides of m + 3. What simplified expression can be used to find the perimeter of the triangle?

OpenStudy (anonymous):

help! any body some good in algebra 1

OpenStudy (anonymous):

can you guys write the steps and how you got the ansewr

OpenStudy (anonymous):

((m+3)*2)+m=P ?

OpenStudy (anonymous):

I'm not sure though

OpenStudy (anonymous):

PERIMETER is sum OF ALL SIDES

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

So if sides are m, and m+3 and m+3 sum is 3m+6

OpenStudy (anonymous):

yea so um idk this is algebra its in equalities i think

OpenStudy (anonymous):

which is your perimeter

OpenStudy (anonymous):

yea so what do i do first

OpenStudy (anonymous):



OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes very sure, do you want to find the perimeter or AREA ?, REMEBER PERIMETER IS ALWAYS the SUM of all sides

OpenStudy (anonymous):

its the perimeter i need to find

OpenStudy (anonymous):

so my formula ((m+3)*2+m)=P works too :)

OpenStudy (anonymous):

yes so thats correct, perimeter is m+ m+3 + m+3 = 3m+6

OpenStudy (anonymous):

i combined the liked terms and i got 3m+6

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

then i divide it by 3 m on both sides which get m=2

OpenStudy (anonymous):

right or wrong?

OpenStudy (anonymous):

why do you want to divide, you just need the perimeter , P= 3m+6 are you supposed to find the value of m, and do you have the value of perimeter given?

OpenStudy (anonymous):

spider, "m" is a variable - it has no number except the one you, or your textbook or teacher give it, if they didn't give you a number for Perimeter variable, or m, then it's just m

OpenStudy (anonymous):

oh okay thanks C: